### v2

@@ Abstract
Gravitational particle production in cosmology is a mechanism through which particles of different natures are produced during the very early universe. It is a general mechanism that explains how the universe became populated with the particles of the Standard Model after cosmic inflation and may also account for the origin of dark matter. In this work, we study the non-perturbative production of massive scalar particles in the Higgs-$R^{2}$ inflation model, a two-field scalar inflation model within the Einstein frame. We consider spectator scalar fields that are conformally or minimally coupled to the gravitational field through the curvature scalar $R$ which in turn is a time-dependent function determined by the fields driving inflationary dynamics. We numerically compute the production of these particles using the Bogolyubov transformation method for each scenario, aiming to assess the spectrum of the produced particles. For both scenarios, we consider light particles with masses $m_{\chi}\ll H_{\mathrm{end}}$ and large masses that exceed the Hubble scale at the end of inflation $m_{\chi}\gtrsim H_{\mathrm{end}}$.
@@ 1 Introduction
Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics.
@@
It describes a period of accelerated expansion driven by a scalar field $\phi$ that slowly rolls towards the minimum of its potential $V(\phi)$. This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2]. During the inflationary epoch, significant particle production occurs, particularly when the kinetic energy of the inflaton field becomes comparable to its potential energy. This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4]. This phenomenon, known as gravitational particle production, is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5]. Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion. Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$. At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance is triggered, leading to instantaneous particle production [6, 7]. However, matter fields need not be directly coupled with the inflaton or standard model particles. Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation. If the contribution of the matter field to the energy density is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry. Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13]. Beyond its role in the origin of dark matter, gravitational particle production is also linked to key early-universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20], and primordial magnetogenesis [21, 22, 23, 24]. Furthermore, it leaves significant observable imprints on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
@@
In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model. This inflation model extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30]. Here, cosmic inflation is driven by the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron. Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31]. At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential minimum. This oscillatory behavior enhances scalar particle production through non-perturbative mechanisms such as tachyonic instabilities [32, 33]. We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that does not contribute to the energy density during the inflationary epoch, with negligible interactions with the SM but considering a direct coupling to the Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$. Initially, the inflationary sector and the spectator field are not explicitly coupled in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field, and free scalars induce couplings upon transforming to the Einstein frame. This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics. Induced coupling in the Einstein frame between free scalars and the scalar sector of the action provides an excellent channel for scalaron decay into light scalars. However, in the non-perturbative approach, the field mass $m_{\chi}$ does not necessarily have to be light, and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, where $H_{\mathrm{end}}$ is the Hubble ratio at the end of inflation. We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$. To achieve this, we employ the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles generated by the background dynamics in Higgs-$R^{2}$ inflation, solving the mode equations for the quantized modes $\chi$.
@@
The article is organized as follows: Section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions. Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, in Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation. We adopt a numerical approach to solve the mode equation for modes that leave 5 e-folds before the horizon crossing during inflation, ensuring that each mode is within the horizon at the start of the numerical simulation. Section 4 examines the production of massive particles during preheating. Finally, Section 5 discusses the results and conclusions.
@@ 2 Brief review of Higgs-$R^{2}$ inflationary model
This expression allows us to make a perturbative treatment of the model. To describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form. So we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$. Thus, to go from the Jordan frame to Einstein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
@@
The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$ and mixes terms which make it more difficult to analyze analytically the background trajectory. Nevertheless, during inflation, we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31] and with the following interesting features
@@
We can see the positions of the valleys and the hill at $h=0$ in figure 1. The inflationary dynamics occur along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation. To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [38] frames. In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
@@ 2.1 Background dynamics
One has to numerically solve the system of equations (16), (17), (18) and (19) to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$. The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (13) inside one of the valleys of the potential. In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$. However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys are an effective single-field description. Then the scalaron field should fulfill the single-field equation
@@
and the evolution of the Higgs field is determined by the condition (13). We can assume a single-field slow roll approximation in the inflationary dynamics inside the valleys. This happens when $\ddot{\phi}\ll 3H\dot{\phi}$, thus the field equation for $\phi$ is
$3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (21)
which can be integrated by quadrature together with the Friedman equations
$\phi(t)=\phi_{0}+\alpha^{-1}\log\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)\,,$ (22)
$a(t)=e^{H_{0}t}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{3/4}\,,$ (23)
$H(t)=H_{0}\,-\dfrac{3}{2}\,t_{\mathrm{end}}^{-1}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{-1}\,,$ (24)
where
$t_{\mathrm{end}}=\dfrac{3}{4}\,H_{0}^{-1}\,e^{\alpha\phi_{0}}\,,\quad H_{0}\simeq\dfrac{M_{p}^{4}}{\sqrt{12}(\xi_{s}+\xi_{h}^{2}/\lambda)}\,,$ (25)
is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation. It is clear from slow-roll solutions that during inflation we have the usual quasi-De Sitter solutions for $\phi$, $a$, and $H$. The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
$\phi(t_{\mathrm{end}})$ $=$ $\alpha^{-1}\log\left(1+\dfrac{2}{\sqrt{3}}\right)\approx 0.92M_{p}\,,$ (26)
$H_{e}$$\approx$ $0.54\,H_{0}\,.$ (27)
@@ 3 Scalar fields in the Einstein frame
In this section, we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initially assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame. In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflationary sector. In the Jordan frame our model is given by
@@
where $\psi_{i}$ are the different matter fields that are present during and after inflation. In this work, we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally. The action of such a field is given by
$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (29)
@@
The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to the gravity sector and may take any value; however, studies based on electroweak vacuum stability [40] suggest that the value of non-minimal coupling constant should not exceed $10$. Moreover there are two interesting special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. The value $\xi_{\chi}=0$ is called the case of minimal coupling. while $\xi_{\chi}=1/6$ corresponds to conformal coupling. The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if, in addition, $m_{\chi}=0$.
@@
Although there is no direct coupling between matter fields and the inflationary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field present during inflation and after inflation, and a gravitational particle production process may occur. In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also occur as long as the matter fields are not conformally invariant, namely, if the scalar field is massive $m_{\chi}\neq 0$. These kinds of models, in which the scalar fields interact only gravitationally, are especially interesting. They can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles. This process of particle creation due to gravitational interactions has implications for the reheating process [17, 41, 42]. It also plays a role in gravitational dark matter production [43, 44, 45, 12, 11, 9, 46].
What follows is to perform the conformal transformation (4) to the action (29) to change to the Einstein frame. Since $\chi$ is a spectator field, it remains energetically subdominant, $\rho_{\chi}\ll 3H^{2}M_{p}^{2}$, throughout the entire background evolution. Therefore, we can apply the same conformal transformation of the metric (4) as in the previous section. By doing so, we avoid considering back-reaction effects on the space-time geometry and the background dynamics. In the Einstein frame, we obtain the following action
@@
where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$. We can perform another conformal transformation which involves the scalar field $\tilde{\chi}$ leading to another Einstein frame to remove the non-minimal coupling constant $\xi_{\chi}$. However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field are responsible for the dynamics during and after inflation. One can see that the conformal transformation (4) induces interaction terms between the scalaron $\phi$, the Higgs field $h$, and the free scalar $\tilde{\chi}$. We can identify the interaction terms from the following action
@@
This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$. We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$). In the paper [8], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively in dark matter particles $\chi$ with masses $m_{\chi}<M/2$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron. For heavy particles, perturbative particle production is kinematically forbidden. However, the production of scalar particles is not limited to perturbative methods; there may be particle production due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch. In our case, the classical time-dependent background is defined by the evolution of the scale factor $a(t)$ and the evolution of the scalar fields $\phi$ and $h$. That is, the matter-free scalar field $\chi$ propagates in a time-dependent classical background, whose dynamics are given by the evolution of the $\phi$ and $h$ fields of the Higgs-$R^{2}$ model. To study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation. Varying the action (30) respect to $\tilde{\chi}$ yields the following equation of motion
@@
Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation. This induces a time-dependent effective mass term, $m_{\mathrm{eff}}^{2}(t),$ for the spectator field $\tilde{\chi}$. Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$. As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the figure 2. By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined. However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (20).

### v3

@@ Abstract
Gravitational particle production in cosmology is a mechanism through which particles of different natures are produced during the very early universe. It is a general mechanism that explains how the universe became populated with the particles of the Standard Model after cosmic inflation and may also account for the origin of dark matter. In this work, we study the non-perturbative production of massive scalar particles in the Higgs-$R^{2}$ inflation model, a two-field scalar inflation model within the Einstein frame. We consider spectator scalar fields that are conformally or minimally coupled to the gravitational field through the curvature scalar $R$ which in turn is a time-dependent function determined by the fields driving inflationary dynamics. We numerically compute the production of these particles using the Bogolyubov transformation method for each scenario, aiming to assess the spectrum of the produced particles. For both scenarios, we consider light particles with masses $m_{\chi}\ll H_{\mathrm{end}}$ and large masses that exceed the Hubble scale at the end of inflation $m_{\chi}\gtrsim H_{\mathrm{end}}$. We use these numerical results to calculate the relic abundance $\Omega_{\chi}h^{2}$ to find out if the model is viable as a dark matter candidate.
@@ 1 Introduction
Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics. It describes a period of accelerated expansion driven by a scalar field $\phi$ that slowly rolls towards the minimum of its potential $V(\phi)$. This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2]. During the inflationary epoch, significant particle production occurs, particularly when the kinetic energy of the inflaton field becomes comparable to its potential energy. This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4]. This phenomenon, known as gravitational particle production, is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5]. Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion. Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$. At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance is triggered, leading to instantaneous particle production [6, 7]. However, matter fields need not be directly coupled with the inflaton or standard model particles. Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation. If the contribution of the matter field to the energy density is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry. Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13]. Beyond its role in the origin of dark matter, gravitational particle production is also linked to key early-universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20], and primordial magnetogenesis [21, 22, 23, 24]. Furthermore, it leaves significant observable imprints on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
@@
In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model. This inflation model extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30]. Here, cosmic inflation is driven by the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron. Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31]. At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential minimum. This oscillatory behavior enhances scalar particle production through non-perturbative mechanisms such as tachyonic instabilities [32, 33]. The preheating phase after inflation has been studied [34, 35, 36, 37], as well as the production of primordial magnetic fields during the inflationary phase [38], considering effective single-field inflation. We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that does not contribute to the energy density during the inflationary epoch, with negligible interactions with the SM but considering a direct coupling to the Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$. Initially, the inflationary sector and the spectator field are not explicitly coupled in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field, and free scalars induce couplings upon transforming to the Einstein frame. This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics. Induced coupling in the Einstein frame between free scalars and the scalar sector of the action provides an excellent channel for scalaron decay into light scalars. However, in the non-perturbative approach, the field mass $m_{\chi}$ does not necessarily have to be light, and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, where $H_{\mathrm{end}}$ is the Hubble ratio at the end of inflation. We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$. To achieve this, we employ the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles generated by the background dynamics in Higgs-$R^{2}$ inflation, solving the mode equations for the quantized modes $\chi$.
@@
The article is organized as follows: Section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions. Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame. Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation. We adopt a numerical approach to solve the mode equation for modes that leave 5 e-folds before the horizon crossing during inflation, ensuring that each mode is within the horizon at the start of the numerical simulation. Section 4 examines the production of massive particles during preheating. In section 5 we show the numerical results for the gravitational production spectra for $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ and for the time-evolution of $\beta_{k}$-Bogolyubov coefficient. Lastly, section 6 is dedicated to the relic abundance of gravitational production of scalars.
@@ 2 Brief review of Higgs-$R^{2}$ inflationary model
This expression allows us to make a perturbative treatment of the model. To describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form. So we need to carry out a conformal transformation $\Psi^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$. Thus, to go from the Jordan frame to Einstein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
@@
The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$ and mixes terms which make it more difficult to analyze analytically the background trajectory.
Nevertheless, during inflation, we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31] and with the following interesting features
@@
We can see the positions of the valleys and the hill at $h=0$ in Fig. 1. The inflationary dynamics occur along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation. To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [43] frames. In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
@@ 2.1 Background dynamics
One has to numerically solve the system of equations (16), (17), (18) and (19) to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$. The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (13) inside one of the valleys of the potential. In the Fig. 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$. However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys are an effective single-field description. Then the scalaron field should fulfill the single-field equation [14, 38]
@@
and the evolution of the Higgs field is determined by the condition (13). It is worth mentioning that the effective single-field approximation is only valid during inflation within one of the valleys of the potential (8). The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final value $\phi(t_{\mathrm{end)}}\approx 0.92M_{p}$, where
$t_{\mathrm{end}}=\dfrac{3}{4}\,H_{0}^{-1}\,e^{\alpha\phi_{0}}\,.$ (21)
After inflation, during the reheating era, the scalaron $\phi$ and the Higgs field $h$ oscillate around the global minimum $(0,0)$ of (8). That is, to follow the complete evolution of the fields after inflation, the system of equations (16), (17), (18), and (19) must be solved numerically with the appropriate initial conditions. Inflation within the valleys occurs for super-Planckian values of the field $\phi\gg M_{p}$ in order to produce sufficient inflation, i.e., $N_{\mathrm{end}}\gtrsim 60$. This condition is satisfied if $\phi_{0}\simeq 5.7M_{p}$, while the value of the Higgs field at the beginning of inflation is obtained by evaluating (13) at $\phi_{0}$.
@@ 3 Scalar fields in the Einstein frame
In this section, we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initially assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame. In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflationary sector. This aims to induce direct couplings between the background composed of the fields $\phi,h$ and the scale factor $a(t)$, while also making it easier to solve the system’s dynamics. In the Jordan frame our model is given by
@@
where $\psi_{i}$ are the different matter fields that are present during and after inflation. In this work, we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally. To the reader interested in the evolution and production of gauge bosons and other kind of matter fields after inflation, we recommend [37] and the references therein. The action in the Jordan frame of a free scalar field $\chi$ is given by
$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}-\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (23)
@@
The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to the gravity sector and may take any value; however, studies based on electroweak vacuum stability [45] suggest that the value of non-minimal coupling constant should not exceed $10$. Moreover there are two interesting special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. The value $\xi_{\chi}=0$ is called the case of minimal coupling while $\xi_{\chi}=1/6$ corresponds to conformal coupling. The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if, in addition, $m_{\chi}=0$.
@@
Although there is no direct coupling between matter fields and the inflationary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field present during inflation and after inflation, and a gravitational particle production process may occur. In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also occur as long as the matter fields are not conformally invariant, namely, if the scalar field is massive $m_{\chi}\neq 0$. These kinds of models, in which the scalar fields interact only gravitationally, are especially interesting. They can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles. This process of particle creation due to gravitational interactions has implications for the reheating process [17, 46, 47], and it also plays a role in gravitational dark matter production [48, 49, 50, 12, 11, 9, 51]. However, as pointed out in [52], the analysis of matter fields with non-minimal coupling in multi-field models of inflation becomes complicated even if the field metric is flat, as moving from one conformal frame to another induces non-canonical kinetic terms. The interpretation of the reheating phase after inflation also changes. While in the Jordan framework it is mainly due to the oscillatory behavior of the Hubble parameter, in the Einstein framework it is mainly due to the explicit interaction terms between the matter fields, and the inflationary sector of the model, through which particle production and reheating of the universe occurs. Despite these differences, evaluating the inflationary dynamics is much simpler in the Einstein frame [53], as in this frame, Einstein’s equations are second-order, making their evaluation more straightforward [54, 55]. Since the computation of the $\beta$-Bogolyubov coefficient is independent of the conformal frame [3], we choose to perform the calculation in the Einstein frame to avoid unnecessary complications when studying the background dynamics in the Jordan frame. While gravitational particle production is typically carried out in the Jordan frame [56, 57, 9, 58], performing the analysis in the Einstein frame [52, 59, 60] is perfectly valid and yields similar results.
As we assume that the energy density of the $\chi$ field never dominates, the reheating phase after inflation cannot proceed purely through gravitational interactions. Instead, it is assumed that reheating occurs via the decay of the scalaron into light particles $m_{\chi}\ll M$ [61, 8, 62] and through the production of gauge bosons due to Higgs boson decay [28, 37, 35]. Thus, the spectator field $\chi$ is considered as dark matter produced gravitationally. Given that the $\chi$-field interacts only gravitationally with the background fields $\phi$ and $h$, its dynamics can be analyzed effectively within the Einstein frame. To study its evolution, we apply the same conformal transformation of the metric (4) as in the previous section. By doing so, we avoid considering back-reaction effects on the space-time geometry and the background dynamics. In the Einstein frame, we obtain the following action:
@@
where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$. We can perform another conformal transformation which involves the scalar field $\tilde{\chi}$ leading to another Einstein frame to remove the non-minimal coupling constant $\xi_{\chi}$. However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field $h$ are responsible for the dynamics during and after inflation. One can see that the conformal transformation (4) induces interaction terms between the scalaron $\phi$, the Higgs field $h$, and the free scalar $\tilde{\chi}$. We can identify the interaction terms from the following action
@@
This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$. We can see that even with $\xi_{\chi}=0$, this interaction does not disappear unless the field $\tilde{\chi}$ is conformally invariant ($m_{\chi}=0\quad\xi_{\chi}=1/6$). Since we are considering $\chi$ as a spectator dark matter field that interacts only gravitationally, it is natural to explore its production mechanisms within inflationary models. In the literature, there are studies on the production of dark matter in the $R^{2}$ model [8, 63, 62, 64], where a scalar field $\chi$ with action (23) is considered. In [8, 63, 62], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively into dark matter particles $\chi$ with masses $m_{\chi}\ll M$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron. The corresponding two-body decay width is given by [61, 8, 63].
$\Gamma_{\phi\to\chi\chi}=\dfrac{1}{192\pi M_{p}^{2}}\left(1-6\xi_{\chi}+2\dfrac{m_{\chi}^{2}}{M^{2}}\right)^{2}\sqrt{1-\dfrac{4m_{\chi}^{2}}{M^{2}}}$ (26)
One can see that for heavy particles $m_{\chi}\gtrsim M$, perturbative particle production is kinematically forbidden. The production of light particles after inflation occurs mainly through this mechanism [8]. However, in [64] the authors study the gravitational production fo heavy dark matter particles with minimal coupling during and after inflation. They find that no additional particle production occurs during inflation if the particle mass is smaller than the Hubble scale $H_{\mathrm{inf}}$ during inflation. However, if the mass is larger, the production after inflation becomes dominant. On the other hand, there have been few attempts to study spectator dark matter field in the Higgs-$R^{2}$ model. In [65], the authors find that the Higgs-$R^{2}$ model is equivalent to a scalar singlet dark matter or Higgs-portal model, constraining the scalar singlet dark matter sector using the dark matter abundance and model parameters. This work, however, does not consider an additional scalar field coupled to the background. In [66], the authors introduce an extra scalar degree of freedom with non-minimal coupling that plays the role of freeze-in dark matter. They also discuss the freeze-in production of dark matter both from the non-thermal scattering during reheating and the thermal scattering after reheating.
In our case, we investigate the gravitational production of the $\chi$ field through non-perturbative mechanisms, taking into account the induced interaction terms in the Einstein frame. Specifically, particle production can occur due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch. The classical time-dependent background in this scenario is determined by the evolution of the scale factor $a(t)$ and the scalar fields $\phi$ and $h$. That is, the spectator field $\chi$, which is initially decoupled from other matter fields, propagates in a time-dependent classical background governed by the evolution of $\phi$ and h in the Higgs-$R^{2}$ model.
To analyze this process, we need to derive the equation of motion for the spectator field $\tilde{\chi}$ and track the evolution of its modes during and after inflation. By varying the action (24) with respect to $\tilde{\chi}$, we obtain the following equation of motion:
@@
Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation. This induces a time-dependent effective mass term, $m_{\mathrm{eff}}^{2}(t),$ for the spectator field $\tilde{\chi}$. Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$. As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the Fig. 2. By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined. However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (20).